OpenStudy (migitmack):

A wooden garden tray with a base area of xy square inches is filled with soil up to a certain height. The tray contains (x2y + 10xy2 + 12xy) cubic inches of soil. The height of the tray is represented by the expression x^2y+10xy^2+12xy/xy Which expression below is the correct simplified form of the height of the tray? 11xy + 12 x + 10y + 12 x + 10xy2 + 12xy x2y + 10xy2 + 12

OpenStudy (mathteacher1729):

Here is a picture which might help you understand what this all means..  Ultimately you are given the volume (it's a polynomial in x and y) and are asked to find the height given the length and width. V = H * L * W We already know L and W (they are x and y) so V = H*XY Solving for H we have H = V/ (XY) The problem is asking you to divide the polynomial x^2y+10xy^2+12xy by xy. OR "factor our xy" from x^2y+10xy^2+12xy Can you factor out the xy from the polynomial x^2y+10xy^2+12xy ?